If (x,1) is a solution to the equation 4x+2y=10 what is the value of x?
Answer by jim_thompson5910(35256)

4x+2y=10

4x+2(1)=10

4x+2=10

4x=10-2

4x=8

x=8/4

x=2

So the solution is x = 2

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------